(a) How many moles of ammonium ions are in 0.753 g of ammonium carbonate? mol

(b) What is the mass, in grams, of 0.0567 mol of iron(III) phosphate? g

(c) What is the mass, in grams, of 4.72 1023 molecules of aspirin, C9H8O4? g

(d) What is the molar mass of a particular compound if 0.050 mol weighs 5.44 g? g/mol
Answer #1

a) (NH4)2CO3 = 96.09 g/mol

moles = 0.753 / 96.09 = 0.0078

moles of ammonium = 2 x 0.0078  

moles of ammonium ion = 0.0156

b)

FePO4 = 150.82 g/mol

mass = moles x molar mass

mass = 0.0567 x 150.82

mass = 8.55 g

c)

C9H8O4 = 180.16 g/mol

6.023 x 1023 molecules = 180.16 g

4.72 x 1023 molecules = 4.72 x 1023 x 180.16 / 6.023 x 1023 = 141.18 g

mass = 141.18 g

d) moles = mass / molar mass

molar mass = mass / moles = 5..44 / 0.050

molar mass = 136 g/mol
Answer #2

Answers:

(a) Moles of ammonium ions in 0.753 g of ammonium carbonate: 


Answer: 0.0157 mol


Calculation:

  1. Molar mass of ammonium carbonate, (NH₄)₂CO₃:

    • N: 2 × 14.01 = 28.02 g/mol

    • H: 8 × 1.01 = 8.08 g/mol

    • C: 1 × 12.01 = 12.01 g/mol

    • O: 3 × 16.00 = 48.00 g/mol


    • Total = 28.02 + 8.08 + 12.01 + 48.00 = 96.11 g/mol

  2. Moles of (NH₄)₂CO₃ in 0.753 g:

    Moles=MassMolar Mass=0.75396.11=0.00783 mol

  3. Each (NH₄)₂CO₃ has 2 NH₄⁺ ions, so:

    Moles of NH₄⁺=0.00783×2=0.0157 mol


(b) Mass of 0.0567 mol of iron(III) phosphate:
Answer: 8.53 g

Calculation:

  1. Formula: FePO₄

  2. Molar mass:

    • Fe: 55.85 g/mol

    • P: 30.97 g/mol

    • O: 4 × 16.00 = 64.00 g/mol

    • Total = 55.85 + 30.97 + 64.00 = 150.82 g/mol

  3. Mass = Moles × Molar Mass:

    0.0567×150.82=8.53 g


(c) Mass of 4.72 × 10²³ molecules of aspirin (C₉H₈O₄):
Answer: 141.26 g

Calculation:

  1. Molar mass of aspirin:

    • C: 9 × 12.01 = 108.09 g/mol

    • H: 8 × 1.01 = 8.08 g/mol

    • O: 4 × 16.00 = 64.00 g/mol

    • Total = 108.09 + 8.08 + 64.00 = 180.17 g/mol

  2. Moles from molecules (Avogadro’s number = 6.022 × 10²³):

    Moles=4.72×10236.022×1023=0.7838 mol

  3. Mass = Moles × Molar Mass:

    0.7838×180.17=141.26 g


(d) Molar mass of a compound (0.050 mol weighs 5.44 g):
Answer: 108.8 g/mol

Calculation:

Molar Mass=MassMoles=5.440.050=108.8 g/mol


answered by: Harshwardhan kunal